# Topology Learning for Heterogeneous Decentralized Federated Learning over Unreliable D2D Networks

### 21 Dec 2023 v1

@@ I Introduction
With the growth of mobile devices in wireless networks, federated learning (FL) has emerged as a promising solution of distributed edge learning, as it ensures user privacy [1, 2]. This paper specifically focuses on the decentralized federated learning (DFL) system deployed on device-to-device (D2D) networks. In this system, devices solely exchange parameters with neighboring nodes based on a specific communication topology to facilitate collaborative training, eliminating the need for a central server. Therefore, DFL mitigates the risk of system vulnerability to failures caused by communication bottlenecks at the server [3, 4].
Figure 1: An illustration of DFL systems deployed in wireless D2D networks.
However, DFL encounters several challenges in the training of models. One key issue in DFL is to tackle with data heterogeneity since edge devices often collect data from distinct areas, leading to serious variations in data distributions [5, 6]. Another serious challenge in DFL is the transmission outages caused by the User Datagram Protocol (UDP) widely used in D2D networks. Usually, DFL deployed in D2D networks faces the bottleneck of limited communication resources [7], often relying on the lightweight UDP [3, 4]. However, the use of UDP compromises reliability, resulting in transmission outages and package errors in inter-device communication [3, 4, 8]. Consequently, parameters shared by DFL clients are affected by both data heterogeneity and unreliable D2D links, further deteriorating the convergence and performance of DFL.
Numerous studies have been dedicated to addressing these issues. Some have proposed device scheduling methods [1, 2, 5] to alleviate the impact of data heterogeneity. However, these approaches are tailored to CFL with reliable links. In a study outlined in [3], the DFL system deployed in unreliable D2D networks was optimized by adjusting mixing weights to enhance performance. Nevertheless, this approach overlooked the influence of data heterogeneity. Additionally, the authors of [6] optimized the topology based on their proposed quantity termed neighborhood heterogeneity to enhance the training of DFL models. However, their focus was limited to scenarios with label distribution shift and the deployment of DFL in ideal D2D networks without unreliable links. In summary, existing research has overlooked the theoretical convergence analysis of DFL considering both arbitrary data heterogeneity and unreliable D2D links, thus neglecting the enhancement of DFL training based on these crucial factors.
@@ II-B Transmission model of unreliable D2D networks
In the subsequent analysis, we examine the aforementioned successful transmission probability and the one-round latency for DFL based on the transmission model of the considered D2D network. For our analysis, we assume orthogonal resource allocation among D2D links [9], and thus define the Signal-to-Noise Ratio (SNR) between two connected D2D clients as $\mbox{SNR}_{i,j}(t):=(P_{tx}h^{i,j}_{t}|d^{i,j}_{t}|^{-2})/\sigma^{2}$. Here, $P_{tx}$ is the transmit power and $\sigma^{2}$ is the power of noise. $h^{i,j}_{t}$ denotes the Rayleigh fading, modeled as an independent random variable $h^{i,j}_{t}\sim\exp(1)$ across links and rounds [4]. $|d^{i,j}_{t}|$ indicates the distance between client $i$ and client $j$. Based on this modeling, we define the successful transmission probability as:
@@
In this paper, we focus on the synchronous implementation of DFL and define the transmission latency in one round as:
@@ III Convergence Analysis
Notice that $\bar{H}$ can be further bounded by $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|\frac{1}{N}\sum_{j=1}^{N}\nabla F_{j}(\mathbf{w}^{j}_{t})-\sum_{j=1}^{N}\theta_{t}^{i,j}\nabla F_{j}(\mathbf{w}^{j}_{t})\|^{2}+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|\sum_{j=1}^{N}\big{(}\theta_{t}^{i,j}\nabla F_{j}(\mathbf{w}^{j}_{t})\odot\mathbf{m}_{t}^{i,j}-\theta_{t}^{i,j}\nabla F_{j}(\mathbf{w}^{j}_{t})\big{)}\|^{2}$. The first term measures the neighborhood discrepancy by disregarding the influence of unreliable links, which is commonly assumed to be bounded [6]. Hence, our paper further focuses on scenarios where the second term, which captures the effect of transmission outages on aggregated gradients, remains bounded to support this assumption.
@@ IV Topology Learning
With the modeling of Gaussian distributions, we can thus analytically calculate $\hat{H}^{k}_{t}(\boldsymbol{\Theta})$ using the below expression:
@@
$\min_{\boldsymbol{\Theta}\in\mathcal{S}}\Big{\{}g(\boldsymbol{\Theta}):=\sum_{k=1}^{m}\hat{H}^{k}_{t}(\boldsymbol{\Theta})+\frac{d\lambda}{N}\sum_{i,j=1}^{N}(\theta^{i,j})^{2}p_{t}^{i,j}(1-p_{t}^{i,j})\Big{\}},$ (10)
@@
Input: Initialization $\boldsymbol{\Theta}_{0}=\mathbf{I}$ and hyperparameters $\lambda\geq 0,K<T,K\in\mathbb{N}^{+}$.
@@ 3             $\boldsymbol{\Theta}_{t}=\boldsymbol{\Theta}_{t-1}$;
6                   $\boldsymbol{\hat{\Theta}}\leftarrow\mbox{Frank-Wolfe}\big{(}g(\boldsymbol{\Theta}_{t},(\boldsymbol{\mu}_{t},\boldsymbol{\sigma}_{t}),\mathbf{P}_{t},\lambda)\big{)}$;
7                   $\boldsymbol{\Theta}_{t}=\boldsymbol{\hat{\Theta}}$;
@@ V Numerical Results
We utilize CNN as the backbone model. Clients perform SGD with a mini-batch size of $128$ and a learning rate of $0.1$. We employ two widely used variants of the CIFAR-10 dataset: Dirichlet CIFAR-10 and Rotated CIFAR-10 [12]. These two variants are chosen to model label skew and feature skew settings respectively. For Dirichlet CIFAR-10, we set the parameter of Dirichlet distribution to $0.1$. For Rotated CIFAR-10, we rotate images clockwise by $3.6^{\circ}$ for each time and distribute them to $100$ clients one by one. We test the average model $\bar{\mathbf{w}}_{t}$ on the global test set following the same distribution as that of all the clients. We compare our method with STL-FW proposed in [6], which only leverages the label distribution shift to learn the topology. Besides, we set the random $r$-regular graph [6] and fully-connected graph as baselines, with uniform weights for all activated links. The hyperparameter $\lambda$ and $K$ of ToLRDUL are set to $0.001$ and $3$ respectively.
@@
Then we investigate the effect of the degree $r$ of nodes on two variants of CIFAR-10. The results are summarized in Table I and Table II. We find that increasing the value of $r$ leads to an enlargement of the transmission latency $l_{t}$ since activating more links may result in a higher number of unreliable links. Denser topologies result in larger values of $l_{t}$, aligning with our intuition. Besides, an excess of activated links in fully-connected graph cause severe package errors and thus degrade the test accuracy. Notice that ToLRDUL achieves the lowest latency and the highest test accuracy under all three settings since it utilizes both data heterogeneity and channel conditions to learn the topology for performing DFL better.

### 11 Mar 2024 v2

@@ I Introduction
With the growth of mobile devices and edge computing techniques in wireless networks, and the rapid development of deep learning technology in mobile communication [1, 2, 3], federated learning (FL) has emerged as a promising solution of distributed edge learning, as it ensures user privacy [4, 5]. This paper specifically focuses on the decentralized federated learning (DFL) deployed on device-to-device (D2D) networks. In this system, devices solely exchange parameters with neighboring nodes based on a specific communication topology to facilitate collaborative training, eliminating the need for a central server. Therefore, DFL mitigates the risk of system vulnerability to failures caused by communication bottlenecks at the server [6, 7].
Figure 1: An illustration of DFL systems deployed in wireless D2D networks. In this system, the sensors serve as the DFL clients. The UDP protocol is used in D2D communication and the D2D links are unreliable.
However, DFL encounters several challenges in the training of models. One key issue in DFL is to tackle with data heterogeneity since edge devices often need to collect data from distinct areas to carry out intelligent tasks [8, 9, 10], which leads to serious variations in data distributions [11, 12, 13]. Another serious challenge in DFL is the transmission outages caused by the User Datagram Protocol (UDP) widely used in D2D networks. Usually, DFL deployed in D2D networks faces the bottleneck of limited communication resources [14], often relying on the lightweight UDP [6, 7]. However, the use of UDP compromises reliability, resulting in transmission outages and package errors in inter-device communication [6, 7, 15]. Consequently, parameters shared by DFL clients are affected by both data heterogeneity and unreliable D2D links, further deteriorating the convergence and performance of DFL.
Numerous studies have been dedicated to addressing these issues. Some have proposed device scheduling methods [4, 5, 11, 16, 17] to alleviate the impact of data heterogeneity. However, these approaches are tailored to CFL with reliable links. In a study outlined in [6], the DFL system deployed in unreliable D2D networks was optimized by adjusting mixing weights to enhance performance. Nevertheless, this approach overlooked the influence of data heterogeneity. Additionally, the authors of [12] optimized the topology based on their proposed quantity termed neighborhood heterogeneity to enhance the training of DFL models. However, their focus was limited to scenarios with label distribution shift and the deployment of DFL in ideal D2D networks without unreliable links. Additionally, the authors in [18, 19] focus on the convergence time optimization in FL for improving the transmission efficiency, but they also ignore both the unreliable links and data heterogeneity. In summary, existing research has overlooked the theoretical convergence analysis of DFL considering both arbitrary data heterogeneity and unreliable D2D links, thus neglecting the enhancement of DFL training based on these crucial factors.
@@ II-B Transmission model of unreliable D2D networks
In the subsequent analysis, we examine the aforementioned successful transmission probability and the one-round latency for DFL based on the transmission model of the considered D2D network. As shown in Fig. 1, we assume that the sensors collecting environmental data serve as the DFL clients and thus we only consider the terrestrial network and static channel in this paper. For our analysis, we assume orthogonal resource allocation among D2D links [22], and thus define the Signal-to-Noise Ratio (SNR) between two connected D2D clients as $\mbox{SNR}_{i,j}(t):=(P_{tx}h^{i,j}_{t}|d^{i,j}_{t}|^{-2})/\sigma^{2}$. Here, $P_{tx}$ is the transmit power and $\sigma^{2}$ is the power of noise. $h^{i,j}_{t}$ denotes the Rayleigh fading, modeled as an independent random variable $h^{i,j}_{t}\sim\exp(1)$ across links and rounds [7]. $|d^{i,j}_{t}|$ indicates the distance between client $i$ and client $j$. Based on this modeling, we define the successful transmission probability as:
@@
In this paper, we focus on the synchronous implementation of DFL. Following the synchronous implementation considered in [7, 23], we assume that D2D devices exchange signaling with the base station providing D2D links for edge devices to achieve consensus. Hence, the transmission latency in one round is defined as:
@@ III Convergence Analysis
Notice that $\bar{H}$ can be further bounded by $\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|\frac{1}{N}\sum_{j=1}^{N}\nabla F_{j}(\mathbf{w}^{j}_{t})-\sum_{j=1}^{N}\theta_{t}^{i,j}\nabla F_{j}(\mathbf{w}^{j}_{t})\|^{2}+\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|\sum_{j=1}^{N}\big{(}\theta_{t}^{i,j}\nabla F_{j}(\mathbf{w}^{j}_{t})\odot\mathbf{m}_{t}^{i,j}-\theta_{t}^{i,j}\nabla F_{j}(\mathbf{w}^{j}_{t})\big{)}\|^{2}$. The first term measures the neighborhood discrepancy by disregarding the influence of unreliable links, which is commonly assumed to be bounded [12]. Hence, our paper further focuses on scenarios where the second term capturing the effect of outages remains bounded to support this assumption.
Consider a scenario where sensors collecting environmental data function as DFL clients to train models for spatially related intelligent tasks, such as indoor localization. However, these sensors are located far apart, resulting in varying spatial distributions of their data and poorer transmission quality. In short, this assumption implies that the impact of data heterogeneity and unreliable links on the aggregated gradients is governed by $\tau$, which reflects the extent of the deviation between the aggregated gradient and the oracle.
@@ IV Topology Learning
Remark 3
Theorem 2 reflects that the proposed unreliable links-aware neighbor- hood discrepancy $\bar{H}$ is impacted by the effect on the aggregated gradients from the expectation of transmission outage, and the variance of this outage.
@@
The rationale for employing the Gaussian distribution as the representation distribution is that the relative entropy in Eq. (8) can be can be computed analytically and we thus avoid using Monte Carlo sampling methods for compute the term in Eq. (8) under other complicated distributions [24]. Therefore, we can analytically calculate $\hat{H}^{k}_{t}(\boldsymbol{\Theta})$ using the below expression:
@@
$\min_{\boldsymbol{\Theta}\in\mathcal{S}}\Big{\{}g_{t}(\boldsymbol{\Theta}):=\sum_{k=1}^{m}\hat{H}^{k}_{t}(\boldsymbol{\Theta})+\frac{d\lambda}{N}\sum_{i,j=1}^{N}(\theta^{i,j})^{2}p_{t}^{i,j}(1-p_{t}^{i,j})\Big{\}},$ (10)
@@
Input: Initialization $\boldsymbol{\Theta}_{0}=\mathbf{I}$ and hyperparameters $\lambda\geq 0,0\leq\gamma\leq 1,1\leq K<T,K\in\mathbb{N}^{+}$.
@@ 3             $\boldsymbol{\Theta}_{t}=\boldsymbol{\Theta}_{t-1}$;
6                   $\tilde{\boldsymbol{\Theta}}=\arg\min_{\mathbf{S}\in\mathcal{S}}<\mathbf{S},\nabla g_{t}(\boldsymbol{\Theta}_{t-1})>$;
7                   $\boldsymbol{\Theta}_{t}=(1-\gamma)\boldsymbol{\Theta}_{t-1}+\gamma\tilde{\boldsymbol{\Theta}}$;
@@ V Numerical Results
We utilize CNN as the backbone model. Clients perform SGD with a mini-batch size of $128$ and a learning rate of $0.1$. The local epoch is set to $1$ and the learning rate decay is set to $0.996$. We employ two widely used variants of the CIFAR-10 dataset: Dirichlet CIFAR-10 and Rotated CIFAR-10 [26]. These two variants are chosen to model label skew and feature skew settings respectively. For Dirichlet CIFAR-10, we set the parameter of Dirichlet distribution to $0.1$. For Rotated CIFAR-10, we rotate images clockwise by $3.6^{\circ}$ for each time and distribute them to $100$ clients one by one. We test the average model $\bar{\mathbf{w}}_{t}$ on the global test set following the same distribution as that of all the clients. In addition, we calculate the global training loss by averaging the local training losses from different clients, and regard it as the metric for convergence. We compare our method with STL-FW proposed in [12], which only leverages the label distribution shift to learn the topology. Specifically, STL-FW requires clients to collect the class proportions $\pi_{ik}=P_{i}(Y=k)$ for each client $i$ and each class $k$ and use Frank-Wolfe method to solve the optimization problem $\min_{\boldsymbol{\Theta}\in\mathcal{S}}\{\frac{1}{n}\|\boldsymbol{\Theta}\Pi-\frac{\mathbf{1}\mathbf{1}^{T}}{n}\Pi\|_{F}^{2}+\frac{\lambda}{n}\|\boldsymbol{\Theta}-\frac{\mathbf{1}\mathbf{1}^{T}}{n}\|_{F}^{2}\}$, where $\Pi\in[0,1]^{n\times K}$ contains the class proportions $\{\pi_{ik}\}$. Besides, we set the random $r$-regular graph [12] and fully-connected graph as baselines, with uniform weights for all activated links. The hyperparameter $\lambda$ and $K$ of ToLRDUL are set to $0.001$ and $3$ respectively.
@@
Then we investigate the effect of the degree $r$ of nodes on two variants of CIFAR-10. The results are summarized in Table I and Table II, where the unit of transmission delay is seconds. We find that increasing the value of $r$ leads to an enlargement of the transmission latency $l_{t}$ since activating more links may result in a higher number of unreliable links. Denser topologies result in larger values of $l_{t}$, aligning with our intuition. Besides, an excess of activated links in fully-connected graph cause severe package errors and thus degrade the test accuracy. Notice that ToLRDUL achieves the lowest latency and the highest test accuracy under all three settings since it utilizes both data heterogeneity and channel conditions to learn the topology for performing DFL better.